OpenStudy (anonymous):

Find a ⋅ b. a = 10i + 9j, b = 4i + 3j

OpenStudy (anonymous):

@Hero @hba

OpenStudy (anonymous):

in vector multiplicatoion you have to multiply i vector with i and j with j a.b=(10i+9j).(4i+3j) =>(10*4)+(9*3)=40+27=67 coz i.i=1 and j.j=1

OpenStudy (anonymous):

Thank you @zaibali.qasmi I really appreciate it! :D

OpenStudy (anonymous):

you are welcome...

OpenStudy (anonymous):

Do you think you could help me with one more if I post it?

OpenStudy (anonymous):

It's just slightly different